VIBRONIC ANALYSIS OF THE B ^{2}A′ - X ^{2}A′′ LASER-INDUCED FLUORESCENCE OF JET-COOLED ETHYLTHIO (C_{2}H_{5}S)

\underline{YUAN-PERN LEE} , MIN-YI SHEN, AND WEN-CHING HUNG , Department of Chemistry, National Tsing Hua University, Hsinchu, TAIWAN 30043 .

The ethylthio (C_{2}H_{5}S) radical was formed by 248-nm laser photolysis of diethyl disulfide (C_{2}H_{5}SSC_{2}H_{5}) or ethyl mercaptan (C_{2}H_{5}SH) in a free jet expansion. The B - X fluorescence excitation spectrum was recorded in the wavelength region 398 - 440 nm. Preliminary data show that the origin lies at 23519.6 cm^{−1} with a main progression of C-S stretching (420.5 cm^{−1}). Other vibrational modes at 256.0, 861.0, 1054.6, and 1203.3 cm^{−1} were also observed for the B state. Detection of hot bands in the excitation spectrum enables assignments of three low-lying modes of the X state: 296.2, 479.9, and 672.8 cm^{−1}. Dispersed fluorescence spectra in the spectral region 410 - 520 nm revealed a main progression of C-S stretching (673 cm^{−1} with additional vibrational modes of the X state near 308, 475, 882, 1076, and 1257 cm^{−1}. Observed vibrational wavenumbers of both B and X states agree with theoretical calculations by the B3LYP/cc-PVDZ+f method (for the X state) and the MP2/DZP method (for both states).